How many pints of a 9% cleaning solution must be mixed with 10 pints of a 13%
cleaning solution to give a 11% solution?

Answer by ptaylor(2198)

Let x=amount of 9% solution needed
Now we know that the amount of pure cleaning solution in the 9% mixture (0.09x) plus the amount of pure cleaning solution in the 13% mixture (0.13*10) has to equal the amount of pure cleaning solution in the final mixture (0.11(10+x)). So our equation to solve is:
0.09x+0.13*10=0.11(10+x) get rid of parens and simplify
0.09x+1.3=1.1+0.11x subtract 1.1 and also 0.09x from each side
0.09x-0.09x+1.3-1.1=1.1-1.1+0.11x-0.09x collect like terms
0.2=0.02x divide each side by 0.02
x=10 pints----------------------------------amount of 9% solution needed
CK
0.09*10+0.13*10=0.11*20
0.9+1.3=2.2
2.2=2.2

Hope this helps---ptaylor